OpenStudy (asdfghj8):

What is the value of (a + b)2 when a = 2/3 and b = -2? I know once you plug it in it looks like: (2/3 + -2)^2 I am not sure how to solve after this, please help.

jimthompson5910 (jim_thompson5910):

think of -2 as -2/1 so you really have 2/3 - 2/1 how can you get the 2/1 to have a denominator of 3?

OpenStudy (asdfghj8):

Multiply top and bottom by 3?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so -2/1 turns into -6/3

jimthompson5910 (jim_thompson5910):

`2/3 - 2/1` becomes `2/3 - 6/3` what comes next?

OpenStudy (asdfghj8):

Subtract 2 by 6 giving -4. Then (-4)^2 would be -16??

jimthompson5910 (jim_thompson5910):

2/3 - 6/3 = (2-6)/3 = -4/3 so we now have (-4/3)^2 don't forget the 3 down below

OpenStudy (asdfghj8):

Oh right, so what's next?

jimthompson5910 (jim_thompson5910):

(-4/3)^2 = (-4/3)*(-4/3) = ???

jimthompson5910 (jim_thompson5910):

up top we have -4 times -4 = 16 down below we have 3 times 3 = 9

jimthompson5910 (jim_thompson5910):

(-4/3)*(-4/3) = 16/9

OpenStudy (asdfghj8):

Ohhhhh! That makes sense now, thank you so much!!

jimthompson5910 (jim_thompson5910):

no problem